# Sparse Autoencoders for Hypothesis Generation

### v1

@@ 4 Methods
Let $e_{i}$ denote the text embedding of $x_{i}$. We train a $k$-sparse autoencoder (Makhzani & Frey, 2014). $k$-sparse autoencoders have successfully generated interpretable features both when applied to intermediate layers of language models, as well as text embeddings (Gao et al., 2024; O’Neill et al., 2024). The SAE encodes a text embedding $e_{i}$ as follows (we use OpenAI’s text-3-embedding-small in our experiments):
$z_{i}$ $=\operatorname{TopK}(W_{\operatorname{enc}}e_{i}+b_{\operatorname{enc}})$ (5)
$\hat{e_{i}}$ $=W_{\operatorname{dec}}z+b_{\operatorname{dec}},$ (6)
where $W_{\operatorname{enc}}\in\mathbb{R}^{M\times D},b_{\operatorname{enc}}\in\mathbb{R}^{M},W_{\operatorname{dec}}\in\mathbb{R}^{D\times M},b_{\operatorname{dec}}\in\mathbb{R}^{D},$ and $\operatorname{TopK}$ sets all activations except the top $k$ to zero. In a basic $k$-sparse autoencoder, the loss for a single input is simply
@@
We follow (Gao et al., 2024; O’Neill et al., 2024) in further adding an auxillary loss to avoid “dead latents.” We describe details and hyperparameters in Appendix B.
@@ 5.3 Baselines
We compare HypotheSAEs against recent, state-of-the-art methods (§2). For fair comparison, we run all methods with the same embeddings and LLMs as our method, unless stated otherwise: OpenAI’s text-3-embedding-small for embeddings; GPT-4o for hypothesis proposal; and GPT-4o-mini for hypothesis annotation. We run each method to produce $H$ hypotheses and evaluate all methods identically.
@@ Acknowledgements
We thank Rishi Jha and Hal Triedman for helpful discussion about sparse autoencoders, Anna Lyubarskaja for helpful discussion about Proposition 3.1, Ruiqi Zhong for helpful discussion about NLParam, Serina Chang for useful comments, Sammi Cheung for writing feedback, and members of the Pierson and Garg groups for comments and discussions.
@@ Appendix A Data Preprocessing
We also download Bills from Pham et al. (2024); the dataset was originally assembled by Hoyle et al. (2022). The bills are from the 110th-114th U.S. Congresses (Jan 2007 - Jan 2017), and the texts and metadata are extracted from GovTrack [^5] https://www.govtrack.us/congress/bills/. Each bill has a topic and subtopic. After filtering out very short bill texts (which may not have been properly transcribed), duplicate bill texts, bills with no subtopic, and bills with infrequent ($<$100) subtopics, there are 20,834 items spanning 70 subtopics. We label the 5,038 bills (24.2%) with the 5 most common subtopics as positives, and all others as negatives.
@@
To preprocess the data, we start by downloading all publicly available data from the Upworthy archive [^6] https://upworthy.natematias.com/index, which includes clicks and impressions for 122,565 distinct headlines (we remove 28K rows from a problematic data range as recently reported by the dataset authors [^7] https://upworthy.natematias.com/2024-06-upworthy-archive-update.html). We then group headlines which have the same test_id meaning that those headlines were randomized against one another. There are 14,128 such groups with at least two headlines. Following Zhou et al. (2024), we take the max click-rate and min click-rate headline in each group, and design a pairwise binary classification task to predict which headline in each pair received higher engagement. That is, we construct tuples (headline A, headline B, label), where headline A and headline B are randomly shuffled, and the label is 1 if headline A received higher click-rate than headline B.
@@
We download the Yelp Open Dataset [^8] https://business.yelp.com/data/resources/open-dataset/, and we filter down to the 4.72M reviews where the business belongs to the ‘Restaurant’ category. From this, we randomly sample 200K reviews for training, 10K reviews for validation, and 10K reviews for heldout evaluation.
@@
We download all congressional speeches from the 109th U.S. Congress, which met from January 2005 to January 2007. These speeches were extracted and released by Gentzkow & Shapiro (2010). We filter speeches which are very short (less than 250 characters) and include only speeches from Republican or Democrat speakers; the dataset is roughly balanced by party (51.7% Republican). Some speeches are very long, so we chunk speeches into ten-sentence excerpts; if a speech has more than five such chunks, we include only the first five (to avoid over-representation of idiosyncratic speeches). Of the remaining chunks, we use 114K for training, 16K for validation, and 12K for heldout eval. If a speech has multiple excerpts, we ensure that all of its excerpts are included in the same split, to avoid cross-split leakage.
@@ Appendix B Hyperparameters
B.1 SAE Training
To train SAEs, we take all hyperparameters besides $M$ and $k$ from O’Neill et al. (2024). We tune $(M,k)$ on each dataset by maximizing validation performance. Specifically, for each combination of $(M,k)$, we evaluate validation AUC (or $R^{2}$ on Yelp) after training an $L_{1}$-regularized predictor with exactly $H$ nonzero coefficients. We constrain $M$ and $k$ to powers of two to tractably limit the search space.
These parameters, broadly, tune the level of granularity of the concepts learned by each SAE neuron: $M$ is the total number of concepts which can be learned across the entire dataset, and $k$ is the number of concepts which can be used to represent each instance. For intuition, assume that with $(M,k)=(16,4)$ on Yelp, the SAE learns a single neuron that fires when reviews mention price; with $(M,k)=(32,4)$, the SAE may learn separate neurons for reviews which mention high prices vs. low prices; with $(M,k)=(32,8)$, the SAE may learn to represent more niche features altogether, such as the number of exclamation marks. In general, increasing either parameter yields more granular features, at the risk of producing some neurons that are uninterpretable or redundant.
@@ B.1 SAE Training
Other hyperparameters are as follows (identical for all experiments):
  • Auxilliary $k$: $2\cdot k$. This parameter is the number of latents for which gradients are computed. That is, we compute a secondary reconstruction of the input using not just the top $k$ latents but the top $\operatorname{auxK}>k$ latents, and include this auxiliary reconstruction’s error in the loss. This addition helps reduce “dead” latents (neurons which never fire).
  • Auxilliary loss coefficient: $1/32$. This parameter is the weight in the loss for the auxiliary reconstruction. (With this term, the full loss is: $\mathcal{L}=||e-\hat{e}_{\text{topK}}||_{2}^{2}+w_{\text{aux}}||e-\hat{e}_{\text{auxK}}||_{2}^{2}$.)
@@
B.2 Neuron interpretation
@@
Choosing the number of candidate interpretations depends on budget. In Appendix C we find that selecting the interpretation with highest fidelity from a pool of 3 candidates increases the resulting fidelity modestly, yet statistically significantly; we therefore use 3 candidate interpretation for all experiments. However, we also try ablating this step—that is, generating a single candidate interpretation and using it directly, without computing fidelity. We find that on Congress, the resulting hypotheses are of similar quality: 14/20 are significant, with AUC 0.70, compared to 15/20 significant and AUC 0.70 when using 3 candidate interpretations [^10] To illustrate this point, we tested solely on Congress; we expect similarly modest differences on other datasets.. This variant of our method, Sae-NoVal, costs as little as BERTopic (Table 3).
@@ Appendix C Supporting Experiments
C.2 Effects of the interpretation procedure on fidelity.
@@
  • Top-random vs. high-weak sampling strategy: Prior work (e.g. Templeton et al. (2024); Bills et al. (2023)) interprets SAE neurons by prompting with the top-activating texts and random texts. Relative to our baseline (prompting with highly-activating texts, but not necessarily top-activating; and weakly-activating texts instead of random), this alternate approach has no significant effect on F1 score, though the interpretations tend to skew towards higher precision and lower recall (because the top-activating examples are more specific than examples in the top decile). Our interpretations tend to have similar values of precision and recall, so we use them instead; but this is a hyperparameter practitioners can change depending on their needs.
@@ Appendix E Proofs
Figure 5: The prompt we use to evaluate surface similarity between the inferred hypotheses and reference topics. The few-shot examples demonstrate three levels of similarity: equivalent meaning (yes), related concepts (related), and unrelated concepts (no).

### v2

@@ 4 Methods
Let $e_{i}$ denote a $D$-dimensional text embedding of $x_{i}$. We train a $k$-sparse autoencoder (Makhzani & Frey, 2014). $k$-sparse autoencoders have successfully generated interpretable features both when applied to intermediate layers of language models, as well as text embeddings (Gao et al., 2024; O’Neill et al., 2024). The SAE encodes a text embedding $e_{i}$ as follows (we use OpenAI’s text-embedding-3-small, where $D=1536$, in our experiments):
$z_{i}$ $=\operatorname{ReLU}(\operatorname{TopK}(W_{\operatorname{enc}}(e_{i}-b_{\mathrm{pre}})+b_{\operatorname{enc}}))$ (5)
$\hat{e_{i}}$ $=W_{\operatorname{dec}}z_{i}+b_{\operatorname{dec}},$ (6)
where $b_{\mathrm{pre}}\in\mathbb{R}^{D},W_{\operatorname{enc}}\in\mathbb{R}^{M\times D},b_{\operatorname{enc}}\in\mathbb{R}^{M},W_{\operatorname{dec}}\in\mathbb{R}^{D\times M},b_{\operatorname{dec}}\in\mathbb{R}^{D},$ and $\operatorname{TopK}$ sets all activations except the top $k$ to zero. In a basic $k$-sparse autoencoder, the loss on one input is
@@
We follow Gao et al. (2024); O’Neill et al. (2024) in further adding an auxillary loss to avoid “dead latents.” We describe details and hyperparameters in Appendix B.
@@ 5.3 Baselines
We compare HypotheSAEs against recent, state-of-the-art methods (§2). For fair comparison, we run all methods with the same embeddings and LLMs as our method, unless stated otherwise: OpenAI’s text-embedding-3-small for embeddings; GPT-4o for hypothesis proposal; and GPT-4o-mini for hypothesis annotation. We run each method to produce $H$ hypotheses and evaluate all methods identically.
@@ 7 Conclusion
Software and Data
Code is available on GitHub, and can be installed via pip with pip install hypothesaes. A notebook to reproduce experimental results in the paper is available on the repository. All data used in the paper are available on HuggingFace. A demo to visualize all SAE neurons and how well they predict the target for Headlines, Yelp, and Congress is available at https://hypothesaes.org/.
@@ Acknowledgements
We thank Rishi Jha and Hal Triedman for helpful discussion about sparse autoencoders, Anna Lyubarskaja for helpful discussion about Proposition 3.1, Ruiqi Zhong for helpful discussion about NLParam, Neil Movva and OpenAI for LLM inference credits, Serina Chang for useful comments, Sammi Cheung for writing feedback, and members of the Pierson and Garg groups for comments and discussions.
@@ Appendix A Data Preprocessing
We also download Bills from Pham et al. (2024); the dataset was originally assembled by Hoyle et al. (2022) using GovTrack [^5] https://www.govtrack.us/congress/bills/. The bills are from the 110th-114th U.S. Congresses (Jan 2007 - Jan 2017). Each bill has a topic and subtopic. After filtering out very short bill texts (which may not have been properly transcribed), duplicate bill texts, bills with no subtopic, and bills with infrequent ($<$100) subtopics, there are 20,834 items spanning 70 subtopics. We label the 5,038 bills (24.2%) with the 5 most common subtopics as positives, and all others as negatives.
@@
To preprocess the data, we start by downloading all publicly available data from the Upworthy archive [^6] https://upworthy.natematias.com/index, which includes clicks and impressions for 122,565 distinct headlines (we remove 28K rows from a problematic data range as recently reported by the dataset authors [^7] https://upworthy.natematias.com/2024-06-upworthy-archive-update.html). We then group headlines which have the same test_id (signifying that those headlines were randomized against one another). There are 14,128 such groups with at least two headlines. Following Zhou et al. (2024), we take the headlines with max and min click-rates in each group, and design a binary classification task to predict which headline in the pair received higher engagement. That is, we construct tuples (headline A, headline B, label), where headline A and headline B are randomly shuffled, and the label is 1 if headline A received higher click-rate than headline B.
@@
To handle the pairwise design, we train the SAE on the embeddings of all unique headlines in the training set, and then compute the activation matrices $Z_{\text{SAE}}^{A}$ and $Z_{\text{SAE}}^{B}$ for all of the headline A’s and headline B’s respectively. We select predictive neurons by identifying columns in the matrix difference $\Delta Z_{\text{SAE}}:=Z_{\text{SAE}}^{A}-Z_{\text{SAE}}^{B}$ which are predictive of $\mathbf{y}$.
@@
We download the Yelp Open Dataset [^8] https://business.yelp.com/data/resources/open-dataset/, and we filter to the 4.72M reviews where the business contains the ‘Restaurant’ tag. From this, we randomly sample training, validation, and heldout sets (200K, 10K, and 10K respectively).
@@
We download all congressional speech transcripts from the 109th U.S. Congress (Gentzkow & Shapiro, 2010) [^9] https://data.stanford.edu/congress_text, which met from January 2005 to January 2007. We filter speeches which are very short (less than 250 characters) and include only speeches from Republican or Democrat speakers; the dataset is roughly balanced (51.7% Republican). If a speech is longer than ten sentences, we split it into ten sentence chunks, and exclude any chunks beyond the first five (to avoid over-representation of specific speeches). For the heldout set, we restrict to one chunk per speech to ensure that our hypotheses generalize to a diverse corpus. Ultimately, we use 114K speech chunks for training, 16K for validation, and 12K for heldout eval. If a speech has multiple excerpts, we ensure that all of them are included in the same split to avoid leakage.
@@ Appendix B Hyperparameters
B.1 Training sparse autoencoders
We train $k$-sparse autoencoders, where each forward pass masks all neurons besides the top-$k$ activating ones (Makhzani & Frey, 2014; Gao et al., 2024). The most influential hyperparameters, which we tune for each dataset, are $M$ and $k$. We choose $(M,k)$ by maximizing validation performance. Specifically, for each combination of $(M,k)$, we evaluate validation AUC (or $R^{2}$ on Yelp) after fitting an $L_{1}$-regularized predictor with exactly $H\ll M$ nonzero coefficients. We constrain $M$ and $k$ to powers of two to tractably limit the search space.
These parameters, broadly, control the level of granularity of the concepts learned by each SAE neuron: $M$ is the total number of concepts which can be learned across the entire dataset, and $k$ is the number of concepts which can be used to represent each instance. For intuition, assume that with $(M,k)=(16,4)$ on Yelp, the SAE learns a single neuron that fires when reviews mention price; with $(M,k)=(32,4)$, the SAE may learn separate neurons for reviews which mention high prices vs. low prices (described as “feature splitting” in Bricken et al. (2023)); with $(M,k)=(32,8)$, the SAE may learn to represent more niche features altogether, such as the number of exclamation marks. In general, increasing either parameter yields more granular features, at the risk of producing some neurons that are uninterpretable or redundant.
@@ B.1 Training sparse autoencoders
Following prior work, we add several optimizations to improve SAE training. Following best practices based on results from Bricken et al. (2023), we (i) tie the weights of the pre-encoder bias and the decoder bias, $b_{\mathrm{pre}}=b_{\operatorname{dec}}$; (ii) initialize $b_{\mathrm{pre}}$ to the geometric median of the training set; (iii) initialize $W_{\operatorname{dec}}=W_{\operatorname{enc}}^{T}$; and (iv) force all decoder rows to have unit norm.
To minimize the number of “dead” latent neurons in the SAE—neurons which are never selected by the $\operatorname{TopK}$ filter—we use an auxiliary loss (Gao et al., 2024; O’Neill et al., 2024). The auxiliary loss term provides nonzero gradients for up to $k_{\text{aux}}$ neurons which were not active in the last several forward passes. We fit these neurons to the residual reconstruction error left by the $\operatorname{TopK}$ neurons. That is, let $z_{i}^{\operatorname{AuxK}}$ consist of the activations of the top $k_{\text{aux}}$ dead neurons in $z_{i}$; then
$\mathcal{L}_{\text{aux}}=||W_{\operatorname{dec}}z_{i}^{\operatorname{AuxK}}-(e_{i}-\hat{e}_{i})||^{2}_{2},$
and the full loss, then, is $\mathcal{L}=\mathcal{L}_{\operatorname{TopK}}+w_{\text{aux}}\mathcal{L}_{\operatorname{AuxK}}$.
Hyperparameters for the auxiliary loss and other training hyperparameters are given below (identical for all experiments):
  • $\bm{k_{\text{aux}}}$: $2\cdot k$; that is, up to $2\cdot k$ dead neurons are used to fit the $\operatorname{TopK}$ reconstruction error.
  • Dead neuron threshold steps: 256. This is the number of consecutive steps for which a neuron was not selected by $\operatorname{TopK}$ to be considered “dead.”
  • Auxilliary loss coefficient $w_{\text{aux}}$: $1/32$. This parameter gives the weight of the auxiliary reconstruction’s loss.
@@
B.2 Interpreting neurons with LLMs
@@
Choosing the number of candidate interpretations depends on budget. In Appendix C we find that selecting the interpretation with highest fidelity from a pool of 3 candidates increases the resulting fidelity modestly, yet statistically significantly; we therefore use 3 candidate interpretations for all experiments. However, we also ablate this step—that is, generating a single candidate interpretation and using it directly, without computing fidelity. On Congress (we did not test on the other datasets), the resulting hypotheses are of similar quality: 14/20 are significant, with AUC 0.70, compared to 15/20 significant and AUC 0.70 when using 3 candidate interpretations [^11] To illustrate this point, we tested solely on Congress; we expect similarly modest differences on other datasets.. This variant of our method, Sae-NoVal, costs as little as BERTopic (Table 3).
@@ Appendix C Supporting Experiments
C.2 Effects of the neuron interpretation procedure on fidelity.
@@
  • Top-random vs. high-weak sampling strategy: Prior work (e.g. Templeton et al. (2024); Bills et al. (2023)) interprets SAE neurons by prompting with the top-activating texts and random texts. Relative to our baseline (prompting with highly-activating texts, but not necessarily top-activating; and weakly-activating texts instead of random), this alternate approach has no significant effect on F1 score, though the interpretations tend to be more specific: that is, they have higher precision and lower recall. This is because a neuron’s absolute top-activating examples generally share a more specific characteristic than examples sampled randomly from the top decile. Though neither strategy wins clearly in terms of F1, we choose high-weak sampling for our experiments because, empirically, top-random produces hypotheses which can be too specific (thereby lowering predictiveness); this is discussed further in B.2. Our code defaults to top-random for simplicity, but practitioners should treat sampling strategy as a tunable hyperparameter that depends on desired hypothesis granularity.
@@ Appendix F Prompts
Figure 5: The prompt we use to evaluate surface similarity between the inferred hypotheses and reference topics. The prompt is lightly edited from Zhong et al. (2024). The few-shot examples demonstrate three levels of similarity: equivalent meaning (yes), related concepts (related), and unrelated concepts (no).